# Alpamayo-R1: Bridging Reasoning and Action Prediction for Generalizable Autonomous Driving in the Long Tail

### v1

@@ 6.4.2 Data Scaling
Figure 13: Impact of training data scale on open-loop driving performance. All models use the 0.5B architecture with identical hyperparameters and fixed total training steps. Models are evaluated on $\mathcal{D}_{\text{overall}}$ held-out test set. *Early stopped due to overfitting.
@@ 6.4.3 Cosmos-Reason Physical AI Capabilities
LingoQA Benchmark. Tab.˜10 presents zero-shot evaluation results on the LingoQA benchmark (Marcu et al., 2024), which assesses vision-language models on driving scene understanding. Our Cosmos-Reason-7B model achieves 66.2% accuracy, outperforming various VLMs including GPT-4V (59.6%), Qwen2-VL-7B (52.6%), Qwen2.5-VL-7B (62.2%), InternVL3.5-8B (58.6%), and DeepSeek-VL-7B (46.4%). This improvement over the baselines demonstrates that Physical-AI-focused SFT significantly improves scene understanding capabilities for autonomous driving contexts, complementing the benefits of model scaling shown in Fig.˜12.
@@ 6.5 Ablation: Action Modality Injection
We demonstrate the effectiveness of adopting a continuous action representation governed by unicycle dynamics with flow matching in Tab.˜11. Specifically, we compare a baseline model trained to auto-regressively predict 6 discrete trajectory tokens against a model of identical size and training data that decodes trajectories via flow matching. The discrete trajectory tokenizer in the baseline auto-regressive model is pre-trained via VQGAN (Esser et al., 2021), which minimizes the number of output discrete tokens to reduce the auto-regressive decoding latency while maintaining low reconstruction error. During inference, we set $\delta_{t}=0.2$, i.e., 5 steps, in flow matching to reduce latency with negligible performance degradation. As shown in Tab.˜11, leveraging a dynamically governed continuous action space through flow-matching yields substantial improvements in both open-loop and closed-loop metrics, enhancing comfort and achieving faster inference speed.
@@ 6.6 Ablation: Efficient Vision Encoding
As can be seen in Tab.˜12, the triplane-based multi-camera tokenizer from Ivanovic et al. (2025) achieves nearly identical minADE6 values as the baseline, while only adding 6.3M parameters and reducing sensor token counts by $3.6\times$. Flex (Yang et al., 2025) is able to achieve more drastic improvements, with a token compression of up to $20\times$ while only adding 61.6M parameters to the overall driving model and matching the driving quality of the baseline.
@@
Model Added Parameters $\downarrow$ Tokens per Image $\downarrow$ Rel. minADE6 $\downarrow$
@@ 6.7 On-Vehicle Road Tests
Real-Time Inference Performance. A critical requirement for on-vehicle deployment is real-time inference capability. We benchmark AR1 on an NVIDIA RTX 6000 Pro Blackwell platform, achieving an end-to-end inference latency of 99ms, within the real-time requirements for autonomous driving (typically 100ms). Tab.˜13 provides a detailed breakdown of the inference pipeline, comparing our approach against alternative design choices. The prefilling stage processes the visual tokens and route information through the transformer layers to generate the key-value cache, which is then used during both reasoning and trajectory decoding.
@@ 7 Conclusion
Future Work. While our current evaluation focuses on internal datasets and the LingoQA benchmark, we plan to extend our assessment to additional public benchmarks for autonomous driving planning and decision-making. This will provide a more comprehensive understanding of Alpamayo-R1’s capabilities across diverse evaluation protocols and enable direct comparison with other state-of-the-art methods in the community. More broadly, several promising research directions remain open. First, policy structuring: while our flow-matching-based trajectory decoder provides kinematically feasible outputs, exploring hierarchical policy architectures that decompose high-level meta-actions into structured motion primitives could further improve interpretability and efficiency. Second, reasoning on demand: our current architecture generates reasoning traces for every input; future work could investigate adaptive mechanisms that selectively invoke reasoning only for safety-critical or ambiguous scenarios, enabling more efficient inference-time computation allocation similar to recent advances in test-time scaling (Yao et al., 2023; OpenAI, 2024); Third, auxiliary task integration: while AR1 focuses on trajectory prediction and causal reasoning, incorporating complementary self-supervised objectives, such as depth estimation, scene flow prediction, or 3D Gaussian Splatting representations, could improve the visual backbone’s semantic understanding; Fourth, world model integration: our current approach predicts actions from observed states; incorporating learned world models could enable forward simulation and counterfactual reasoning, improving robustness in dynamic scenarios.
Open Source Release. We plan to release Alpamayo-R1 models on Hugging Face, along with a subset of the CoC dataset, augmenting the sensor data and labels available at NVIDIA’s Hugging Face webpage, to advance research at the intersection of language-based reasoning and autonomous driving.

### v2

@@ 6.5.2 Data Scaling
Figure 13: Impact of training data scale on open-loop driving performance. All models use the 0.5B architecture with identical hyperparameters and fixed total training steps. Models are evaluated on $\mathcal{D}_{\text{overall}}$ held-out test set.
@@ 6.5.3 Cosmos-Reason Physical AI Capabilities
LingoQA Benchmark. Tab.˜11 presents zero-shot evaluation results on the LingoQA benchmark (Marcu et al., 2024), which assesses vision-language models on driving scene understanding. Our Cosmos-Reason-7B model achieves 66.2% accuracy, outperforming various VLMs including GPT-4V (59.6%), Qwen2-VL-7B (52.6%), Qwen2.5-VL-7B (62.2%), InternVL3.5-8B (58.6%), and DeepSeek-VL-7B (46.4%). This improvement over the baselines demonstrates that Physical-AI-focused SFT significantly improves scene understanding capabilities for autonomous driving contexts, complementing the benefits of model scaling shown in Fig.˜12.
@@ 6.6 Ablation: Action Modality Injection
We demonstrate the effectiveness of adopting a continuous action representation governed by unicycle dynamics with flow matching in Tab.˜12. Specifically, we compare a baseline model trained to auto-regressively predict 6 discrete trajectory tokens against a model of identical size and training data that decodes trajectories via flow matching. The discrete trajectory tokenizer in the baseline auto-regressive model is pre-trained via VQGAN (Esser et al., 2021), which minimizes the number of output discrete tokens to reduce the auto-regressive decoding latency while maintaining low reconstruction error. During inference, we set $\delta_{t}=0.2$, i.e., 5 steps, in flow matching to reduce latency with negligible performance degradation. As shown in Tab.˜12, leveraging a dynamically governed continuous action space through flow-matching yields substantial improvements in both open-loop and closed-loop metrics, enhancing comfort and achieving faster inference speed.
@@ 6.7 Ablation: Efficient Vision Encoding
As can be seen in Tab.˜13, the triplane-based multi-camera tokenizer from Ivanovic et al. (2025) achieves nearly identical minADE6 values as the baseline, while only adding 6.3M parameters and reducing sensor token counts by $3.6\times$. Flex (Yang et al., 2025) is able to achieve more drastic improvements, with a token compression of up to $20\times$ while only adding 61.6M parameters to the overall driving model and matching the driving quality of the baseline.
@@
Vision Encoder Added Parameters $\downarrow$ Tokens per Image $\downarrow$ Rel. minADE6 $\downarrow$
@@ 6.8 On-Vehicle Road Tests
Real-Time Inference Performance. A critical requirement for on-vehicle deployment is real-time inference capability. We benchmark AR1 on an NVIDIA RTX 6000 Pro Blackwell platform, achieving an end-to-end inference latency of 99ms, within the real-time requirements for autonomous driving (typically 100ms). Tab.˜14 provides a detailed breakdown of the inference pipeline, comparing our approach against alternative design choices. The prefilling stage processes the visual tokens and route information through the transformer layers to generate the key-value cache, which is then used during both reasoning and trajectory decoding.
@@ 7 Conclusion
Future Work. Several promising research directions remain open. First, policy structuring: while our flow-matching-based trajectory decoder provides kinematically feasible outputs, exploring hierarchical policy architectures that decompose high-level meta-actions into structured motion primitives could further improve interpretability and efficiency. Second, reasoning on demand: our current architecture generates reasoning traces for every input; future work could investigate adaptive mechanisms that selectively invoke reasoning only for safety-critical or ambiguous scenarios, enabling more efficient inference-time computation allocation similar to recent advances in test-time scaling (Yao et al., 2023; OpenAI, 2024); Third, auxiliary task integration: while AR1 focuses on trajectory prediction and causal reasoning, incorporating complementary self-supervised objectives, such as depth estimation, scene flow prediction, or 3D Gaussian Splatting representations, could improve the visual backbone’s semantic understanding; Fourth, world model integration: our current approach predicts actions from observed states; incorporating learned world models could enable forward simulation and counterfactual reasoning, improving robustness in dynamic scenarios.
Open Source Release. We release Alpamayo-R1-10B model weights at https://huggingface.co/nvidia/Alpamayo-R1-10B and inference code at https://github.com/NVlabs/alpamayo. The model is evaluated on the PhysicalAI-AV dataset (NVIDIA, 2025b) and the AlpaSim public scenario set (NVIDIA et al., 2025b), enabling reproducible benchmarking by the research community.